OpenStudy (anonymous):

can some one give me the concept of arc and curvature( related to calculus 3)

OpenStudy (anonymous):

curvature is the rate where the unit vector T turns per unit of length. it is calculated \[k=|dT/ds|\]

OpenStudy (anonymous):

if you are given a position, you can find T by dividing the derivative of position over the absolute value of it. it can be looked at as velocity over speed\[r'(t)/|r'(t)|\]

OpenStudy (anonymous):

ds refers to arc length parameter for a curve. if you're given a curve r(t)=x(t)i+y(t)j+z(t)k, \[s(t)= \int\limits_{t _{0}}^{t} (dx/dt)^{2}+(dy/dt)^{2}+(dz/dt)^{2} dt\]

OpenStudy (anonymous):

now, k can be re written as \[k=|(dT/dt)(dt/ds)|\]

OpenStudy (anonymous):

hmm.........ok thanx

OpenStudy (anonymous):

elica?

OpenStudy (anonymous):

the purpose of finding an arc length can be thought of as wanting to know how sharp a curve is turning

OpenStudy (anonymous):

i just finished this class so my explanation might not be detailed enough since i need to understand more but i shared what i can

OpenStudy (anonymous):

in which class u r?

OpenStudy (anonymous):

calc III just had the finals on monday